OpenStudy (anonymous):

Part A: The Sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the Sun produce in 2.45 ⋅ 10^5 seconds? Part B: Which is the more reasonable measurement of the distance between the tracks on a DVD: 7.4 ⋅ 10^−4 mm or 7.4 ⋅ 10^4 mm? Justify your answer.

OpenStudy (anonymous):

help please

OpenStudy (phi):

for part A, you have ergs/second and seconds and you want ergs notice \[ \frac{ergs}{\cancel{sec}} \cdot \cancel{sec} = ergs \] the point is, you multiply the two numbers can you do that ?

OpenStudy (phi):

you can think of it as multiplying 4 numbers do the "ordinary" numbers first then do the base 10 numbers

OpenStudy (phi):

\[ 3.9 \cdot 10^{33} \cdot 2.45 \cdot 10^5 \]

OpenStudy (anonymous):

wait i got the same answer?

OpenStudy (phi):

I just wrote down the multiplication but you can simplify it. when you multiply, you can change the order, right ? so you could write the problem like this \[ 3.9 \cdot 2.45 \cdot 10^{33} \cdot 10^5 \] can you multiply the first pair ?

OpenStudy (anonymous):

i got 9.945 x 10^165

OpenStudy (anonymous):

i dont think thats right though

OpenStudy (anonymous):

@phi ?

OpenStudy (phi):

you don't multiply exponents, you add them

TheSmartOne (thesmartone):

\(\sf 3.9 \times 2.45=9.555\) not 9.45 :P Use this formula to calculate what \(\sf 10^{33}\times 10^5=?\) \(\sf\Large a^b\times a^c=a^{b+c}\)

OpenStudy (anonymous):

9.945 x 10^38?

OpenStudy (anonymous):

oh

TheSmartOne (thesmartone):

9.555* :P

OpenStudy (anonymous):

9.555 x 10^38

OpenStudy (phi):

it might be easier to remember, if you know that the "exponent" idea is short hand if you say \(5^3\) that means 5*5*5 if you multiply \(5^3 \cdot 5^3\) that is 5*5*5 * 5*5*5 you see you get 5 times itself 6 times. notice you get 6 if you add the exponents

OpenStudy (anonymous):

oh now i get it

OpenStudy (anonymous):

ok so then 9.555 x 10^38 is corrrect for part A

TheSmartOne (thesmartone):

I'll let @phi help you with the rest of the question because when multiple people help, the asker gets confused :P

OpenStudy (anonymous):

ok thank you @TheSmartOne

OpenStudy (anonymous):

so then whats the formula for part B?

OpenStudy (anonymous):

i think its the second one

OpenStudy (anonymous):

7.4 x 10^4

OpenStudy (anonymous):

am i correct?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

am i?

OpenStudy (anonymous):

because that would make 74,000mm

OpenStudy (anonymous):

@phi?

OpenStudy (anonymous):

help?

OpenStudy (anonymous):

anyone? i know you are looking at this...it shows whos veiwing it

OpenStudy (anonymous):

@welshfella @jdoe0001